OpenStudy (coolkid18):

What is the probability of getting two heads and one tail when a coin is flipped three times? 1/3 3/8 1/2

OpenStudy (tgstudios):

i would go with 3/8 because there are 3 things you want to flip and each of them has a chance of getting the other 2 one time... if that's confusing i don't blame you, go with 3/8

Vocaloid (vocaloid):

consider each flip separately probability of heads = 1/2 probability of tails = 1/2 so the probability of getting two heads and a tail is probability of heads * probability of heads * probability of tails = ?